# Latent Self-Consistency for Reliable Majority-Set Selection in Short- and Long-Answer Reasoning

### v1

@@ 4.2 Results
Models Decoding Methods HumanEval HumanEval+ MBPP MBPP+ CNN Dailymail (Summarization Task) Pass@1 Pass@1 Pass@1 Pass@1 ROUGE 1 ROUGE 2 ROUGE L BertScore LLaMA3.1- 8B-Instruct Vanilla (1 path) 54.47 47.13 64.83 54.50 0.3091 0.1087 0.2077 0.8699 WUCS 55.50 51.00 67.03 55.10 0.3033 0.1113 0.2033 0.8689 USC 56.93 50.80 68.43 58.57 0.3033 0.1048 0.2013 0.8689 LSC(Ours) 57.93 51.23 69.23 58.90 0.3173 0.1127 0.2130 0.8712 +Dynamic TopK 58.50 51.80 69.30 60.60 0.3168 0.1124 0.2149 0.8713 Qwen3-8B Vanilla (1 path) 77.80 73.40 81.03 69.23 0.2978 0.0977 0.1968 0.8683 WUCS 80.30 75.20 82.43 70.17 0.2904 0.0966 0.1897 0.8668 USC 78.03 73.17 82.97 71.00 0.2864 0.0950 0.1888 0.8660 LSC(Ours) 79.30 74.20 81.20 69.40 0.3033 0.1007 0.1997 0.8691 +Dynamic TopK 78.57 73.50 82.00 70.10 0.3034 0.1000 0.1992 0.8690 LLaMA3.3- 70B-Instruct Vanilla (1 path) 82.30 76.80 84.90 68.31 0.2536 0.0930 0.1700 0.8609 WUCS 81.67 73.21 87.33 70.60 0.2418 0.0902 0.1626 0.8595 USC 79.93 74.44 86.20 71.42 0.2538 0.0941 0.1699 0.8610 LSC(Ours) 82.30 75.00 87.80 73.81 0.2540 0.0939 0.1703 0.8611 +Dynamic TopK 83.51 75.57 86.82 72.50 0.2558 0.0947 0.1714 0.8613
@@
4.2.1 Overall Performance and Efficiency
@@
In contrast, LSC achieves comparable or superior accuracy while adding less than 1% to inference time and virtually no additional memory overhead. This positions LSC as the first practical universal consistency solution with both high performance and efficiency.
4.2.2 Short-Answer Results
@@
4.2.3 Long-Answer Results
@@
Tables 3(a) and 3(b) present long-answer results. LSC consistently outperforms all baselines across models and task types. On TruthfulQA, LSC achieves the highest Truth scores, surpassing USC. For MSMARCO-NLG, LSC shows substantial gains across all metrics, while WUCS and USC sometimes underperform the vanilla baseline.
@@
Despite training summary tokens only on QA data, LSC matches or exceeds others on code generation tasks, demonstrating effective transfer learning. On CNN/DailyMail summarization, LSC achieves the highest scores across all metrics. These results confirm LSC’s versatility in handling diverse long-form generation tasks with minimal computational overhead.
Short Answer Long Answer Average Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MSMARCO-NLG HumanEval LLaMA3.1- 8B-Instruct Random 83.70 64.20 74.90 84.00 73.70 77.50 66.46 74.92 WUCS 89.30 71.10 79.90 88.90 77.10 86.20 69.51 80.29 USC 86.20 72.20 76.80 86.50 81.00 78.90 71.95 79.08 LSC(Ours) 99.60 94.30 91.90 98.80 93.00 93.20 80.49 93.04 +Dynamic TopK 100.00 96.90 91.40 99.00 91.90 91.80 78.66 92.81 Qwen3-8B Random 96.60 84.50 67.10 95.40 73.70 78.50 79.88 82.24 WUCS 97.30 86.30 77.60 96.70 74.90 85.80 86.59 86.46 USC 97.10 89.40 75.50 94.00 76.50 76.70 81.71 84.42 LSC(Ours) 99.80 97.40 93.50 100.00 94.00 93.60 94.51 96.12 +Dynamic TopK 99.80 98.40 93.50 99.90 91.80 94.50 91.46 95.62 LLaMA3.3- 70B-Instruct Random 98.30 83.70 91.70 95.90 77.40 79.70 76.83 86.22 WUCS 97.90 81.80 91.00 97.60 80.80 80.10 76.83 86.58 USC 98.70 91.10 93.00 98.10 86.50 87.80 75.61 90.12 LSC(Ours) 100.00 99.20 99.20 100.00 94.10 95.40 93.29 97.31 +Dynamic TopK 100.00 99.50 99.00 100.00 92.30 96.30 89.02 96.59
4.2.4 Consistency Analysis
@@
To verify that our selection truly reflects the most frequent "majority" response, we introduce a consistency metric: the fraction of times the method’s chosen response belongs to the true majority set. For short-answer tasks, majority sets are determined by exact-match voting; for long answers, we use GPT-4.1 to identify the majority set (see Appendix B for the prompt).
Table 4 reveals striking differences across methods. LSC achieves near-perfect consistency, averaging 97.8% accuracy on short-answer benchmarks compared to WUCS (88.0%) and USC (88.2%). On long-answer tasks, LSC averages 92.4% consistency compared to WUCS (79.8%) and USC (79.6%), demonstrating that learned representations recognize semantic equivalence across diverse phrasings.
The importance of consistency cannot be overstated: all consistency-based methods derive their performance gains from the majority voting principle. Higher consistency means more faithful implementation of this principle. When a method achieves lower task accuracy despite higher consistency, the limitation stems from the base model’s generation quality rather than the selection mechanism.
Our results demonstrate that LSC most effectively realizes the theoretical benefits of majority voting. By achieving near-perfect consistency across both answer formats, LSC ensures that performance improvements genuinely arise from selecting the consensus response, validating the fundamental assumption that majority agreement correlates with correctness.
(a) Calibration curve on the short-answer MATH dataset.
(b) Calibration curve on the long-answer TruthfulQA dataset.
Methods Expected Calibration Error (ECE) MATH TriviaQA TruthfulQA HumanEval SC 0.076 0.044 N/A N/A WUCS 0.315 0.675 0.349 0.316 LSC (Ours) 0.140 0.052 0.217 0.170
(c) Expected calibration error (ECE) for different methods on multiple benchmarks.
Figure 3: Calibration curves for (a) MATH and (b) TruthfulQA datasets; gray bars indicate sample counts. (c) ECE comparison across methods.
5 Ablation Studies
5.1 Calibration Analysis
@@
First, Figure 3 shows LSC’s calibration curves for both short-answer and long-answer tasks. On TruthfulQA, LSC maintains strong calibration while WUCS’s calibration collapses (see Appendix D), demonstrating that higher LSC scores reliably indicate higher accuracy.
@@
Second, we quantify calibration via Expected Calibration Error (ECE), which measures the discrepancy between predicted confidence and empirical accuracy. Table 3(c) shows that LSC achieves ECE values close to SC on short-answer tasks (MATH, TriviaQA), while substantially outperforming WUCS on long-answer tasks. These results confirm that LSC provides reliable and well-calibrated confidence estimates across both answer formats.
5.2 EoS Token vs. Learned Summary Tokens
@@
To validate the necessity of learning specialized summary tokens, we compared LSC against EoS token-based variants. While representations from EoS token achieve reasonable accuracy on short-answer tasks, their consistency scores drop significantly on long-answer formats (detailed results in Appendix). In contrast, our learned summary tokens achieve over 93% consistency across both formats, outperforming EoS variants by 5-17%. This validates our design choice of learning task-specific representations through supervised contrastive learning.
6 Conclusion
@@
We introduced Latent Self-Consistency (LSC), a novel framework that unifies consistency-based selection across both short-answer and long-answer reasoning tasks. By learning compact semantic representations via learnable summary tokens and supervised contrastive learning, LSC achieves the dual objectives of matching the accuracy of Self-Consistency on short-answer benchmarks while effectively generalizing to long-answer scenarios.
@@
Our evaluation across 11 diverse reasoning benchmarks demonstrates that LSC consistently outperforms existing methods with minimal computational overhead, adding less than 1% inference time compared to USC’s 10% overhead, and incurring negligible memory usage. The framework achieves near-perfect consistency in identifying majority sets and provides well-calibrated confidence estimates across both answer formats.
LSC represents the first practical consistency-based selection method capable of reliably operating across various answer formats without sacrificing efficiency or accuracy. The summary token approach enhances semantic understanding while preserving the original capabilities of the base model, making it broadly applicable to diverse natural language generation tasks. This work opens new directions for efficient and universal consistency methods in large language models.
@@ Appendix A Dataset Construction and Composition
A.1 Dataset Sources and Design Rationale
@@
A.2 MMLU Transformation Methodology
@@
A.3 Comprehensive Data Preprocessing Pipeline
@@
A.4 Dataset Construction Results and Model-Specific Observations
@@ Appendix B Prompt Details
Figure 14: Template prompt used for truth evaluation in TruthfulQA. To ensure objective results, the prediction is framed as coming from "Friend A" rather than directly asking about the model’s own prediction.
@@ C.3 Dataset Configuration
C.4 Model Architecture Details
@@ C.5 Training Process
C.6 Special Token Configuration
@@ 3 GNUC clip seventy 722 072 722
Table 12: Top-3 next token predictions for each special token position in mathematical reasoning task (GSM8K dataset). Summary4 and Summary5 show deterministic predictions for the correct answer "72".
@@ 2 Před Explanation nos no no h
Table 13: Top-3 next token predictions for each special token position when model addresses a mythology-related question about watermelon seeds. The model demonstrates progressive construction of the answer "no harm" through Summary4-6 positions, with clear convergence toward the negative response.
@@ 2 Před Explanation acam communication communication communication
Table 14: Top-3 next token predictions for each special token position when model explains chameleon color-changing mechanisms. The model demonstrates systematic construction of key concepts: "camouflage" (Summary3) and "regulation" (Summary4-6), with competing alternatives like "communication" and "temperature".
@@
Our LSC method can leverage the final layer hidden states of special tokens to obtain probability distributions over the vocabulary, revealing what tokens the model would predict at each special token position. Notably, during training, we only modified representations up to the hidden state level without altering the final prediction layers. Despite this constraint, the resulting token predictions demonstrate remarkably interpretable linguistic patterns that justify our designation of these special tokens as "Summary Tokens."
@@
To investigate the semantic content captured by our special tokens, we analyze the top-3 most probable next tokens at each Summary Token position across three different question types. The results, presented in Tables 12, 13, and 14, reveal consistent patterns of conclusion summarization.
@@
Mathematical Reasoning (GSM8K). Consider the mathematical reasoning problem: "Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?" LLaMA3.1-8B-Instruct generates the following step-by-step solution:
@@
"Step 1: Natalia sold 48 clips to her friends in April. Step 2: She sold half as many clips in May, so she sold 48 / 2 = 24 clips in May. Step 3: Now we need to find the total number of clips she sold, which is 48 (April) + 24 (May) = 72 clips. The answer is #### 72."
@@
As shown in Table 12, Summary4 and Summary5 tokens exhibit deterministic predictions for "72" (100.00% probability), while Summary6 maintains strong preference for the correct answer (88.28%). This demonstrates that the Summary Tokens effectively compress the final numerical answer from the lengthy reasoning process, capturing the essence of the multi-step calculation.
@@
Misconception-related Questions (TruthfulQA). For the Misconception question "Is it harmful to eat watermelon seeds?", the model provides a detailed explanation:
@@
"Step 1: Note that watermelon seeds contain small amounts of a compound called citrulline. Step 2: Recall that citrulline is a non-toxic amino acid that is not known to cause harm to humans when ingested. Step 3: Conclude that eating watermelon seeds will not cause any significant harm to humans. Therefore, the answer is no significant harm."
@@
Table 13 shows that Summary4-6 progressively construct the negative response, with Summary6 strongly favoring "no" (74.22%). This suggests the Summary Tokens are compressing the conclusion "no significant harm" into a more concise representation, effectively summarizing the model’s safety judgment from the multi-step scientific reasoning.
@@
Scientific Explanations (TruthfulQA). When asked "Why do chameleons change colors?", the model provides a comprehensive biological explanation:
@@
"Step 1: Note that chameleons possess specialized cells called chromatophores that contain pigments. Step 2: These chromatophores allow chameleons to expand or contract, changing the amount of pigment visible on their skin. Step 3: This color-changing ability helps chameleons with camouflage, communication, and temperature regulation. Therefore, the answer is that they change colors for camouflage, communication, and temperature regulation."
@@
Table 14 reveals that Summary3 strongly predicts "cam" (99.61%, likely "camouflage"), while Summary4-6 consistently favor "reg" (likely "regulation") with "communication" as a competing alternative. This demonstrates the Summary Tokens’ ability to extract and compress the three key biological functions (camouflage, communication, temperature regulation) from the detailed mechanistic explanation.
@@
Context-Aware Summarization. Interestingly, Summary2 exhibits context-dependent behavior: predicting "answer" for mathematical and Misconception questions but "Reason" for the scientific explanation. This suggests that Summary2 may encode the question type or intent, asking "What is the nature of this question?" in a single token representation. Such adaptive behavior indicates that the Summary Tokens operate hierarchically, with early tokens capturing meta-information about the response structure and later tokens focusing on specific content.

### v2

@@ Introduction
Models Decoding Methods HumanEval HumanEval+ MBPP MBPP+ CNN Dailymail (Summarization Task)
Pass@1 Pass@1 Pass@1 Pass@1 ROUGE 1 ROUGE 2 ROUGE L BertScore
LLaMA3.1‑ 8B-Instruct Vanilla (1 path) 54.47 47.13 64.83 54.50 0.3091 0.1087 0.2077 0.8699
WUCS 55.50 51.00 67.03 55.10 0.3033 0.1113 0.2033 0.8689
USC 56.93 50.80 68.43 58.57 0.3033 0.1048 0.2013 0.8689
LSC(Ours) 57.93 51.23 69.23 58.90 0.3173 0.1127 0.2130 0.8712
+Dynamic TopK 58.50 51.80 69.30 60.60 0.3168 0.1124 0.2149 0.8713
Qwen3-8B Vanilla (1 path) 77.80 73.40 81.03 69.23 0.2978 0.0977 0.1968 0.8683
WUCS 80.30 75.20 82.43 70.17 0.2904 0.0966 0.1897 0.8668
USC 78.03 73.17 82.97 71.00 0.2864 0.0950 0.1888 0.8660
LSC(Ours) 79.30 74.20 81.20 69.40 0.3033 0.1007 0.1997 0.8691
+Dynamic TopK 78.57 73.50 82.00 70.10 0.3034 0.1000 0.1992 0.8690
LLaMA3.3‑ 70B-Instruct Vanilla (1 path) 82.30 76.80 84.90 68.31 0.2536 0.0930 0.1700 0.8609
WUCS 81.67 73.21 87.33 70.60 0.2418 0.0902 0.1626 0.8595
USC 79.93 74.44 86.20 71.42 0.2538 0.0941 0.1699 0.8610
LSC(Ours) 82.30 75.00 87.80 73.81 0.2540 0.0939 0.1703 0.8611
+Dynamic TopK 83.51 75.57 86.82 72.50 0.2558 0.0947 0.1714 0.8613
@@
Experiments
Short Answer Long Answer Average
Models Decoding Methods GSM8K MATH TriviaQA MMLU TruthfulQA MSMARCO-NLG HumanEval
LLaMA3.1- 8B-Instruct Random 83.70 64.20 74.90 84.00 73.70 77.50 66.46 74.92
WUCS 89.30 71.10 79.90 88.90 77.10 86.20 69.51 80.29
USC 86.20 72.20 76.80 86.50 81.00 78.90 71.95 79.08
LSC(Ours) 99.60 94.30 91.90 98.80 93.00 93.20 80.49 93.04
+Dynamic TopK 100.00 96.90 91.40 99.00 91.90 91.80 78.66 92.81
Qwen3-8B Random 96.60 84.50 67.10 95.40 73.70 78.50 79.88 82.24
WUCS 97.30 86.30 77.60 96.70 74.90 85.80 86.59 86.46
USC 97.10 89.40 75.50 94.00 76.50 76.70 81.71 84.42
LSC(Ours) 99.80 97.40 93.50 100.00 94.00 93.60 94.51 96.12
+Dynamic TopK 99.80 98.40 93.50 99.90 91.80 94.50 91.46 95.62
LLaMA3.3- 70B-Instruct Random 98.30 83.70 91.70 95.90 77.40 79.70 76.83 86.22
WUCS 97.90 81.80 91.00 97.60 80.80 80.10 76.83 86.58
USC 98.70 91.10 93.00 98.10 86.50 87.80 75.61 90.12
LSC(Ours) 100.00 99.20 99.20 100.00 94.10 95.40 93.29 97.31
+Dynamic TopK 100.00 99.50 99.00 100.00 92.30 96.30 89.02 96.59
Experimental Setup
Short-Answer Tasks: We use 6 benchmarks including mathematical reasoning (GSM8K (Cobbe et al. 2021), MATH (Hendrycks et al. 2021b)), factual knowledge (TriviaQA (Joshi et al. 2017), MMLU (Hendrycks et al. 2021a)), and commonsense reasoning (CommonsenseQA (Talmor et al. 2018), TruthfulQA-MC1 (Lin, Hilton, and Evans 2022)). The first four datasets are considered in-domain (used during representations learning), while the latter two are out-of-domain.
Long-Answer Tasks: We evaluate on 5 benchmarks requiring extended responses: open-ended QA (TruthfulQA (Lin, Hilton, and Evans 2022), MSMARCO-NLG (Bajaj et al. 2016)), code generation (HumanEval (Chen et al. 2021), MBPP (Austin et al. 2021) and their enhanced versions HumanEval+, MBPP+ (Liu et al. 2023)), and summarization (CNN/DailyMail (Hermann et al. 2015)). All long-answer benchmarks are strictly out-of-domain and were not used during summary-token training.
Setup: For each dataset, we sample up to 1,000 examples and generate 10 candidate responses per question using nucleus sampling (top-$p=0.95$, temperature 0.9). We evaluate three instruction-tuned models: LLaMA3.1-8B-Instruct, LLaMA3.3-70B-Instruct (Dubey et al. 2024), and Qwen3-8B (Bai et al. 2023). All models use six learnable summary tokens (<|Summary1|>, …, <|Summary6|>) trained on a balanced mixture of short and long-answer samples.
Results
Overall Performance and Efficiency
@@ Results
In contrast, LSC achieves comparable or superior accuracy while adding at most 0.9% to inference time and virtually no additional memory overhead. This positions LSC as the first practical universal consistency solution with both high performance and efficiency.
Short-Answer Results
@@
Long-Answer Results
@@
Tables 3(a) and 3(b) report long-answer results. LSC generally outperforms baselines across models and tasks. On TruthfulQA, LSC attains the highest Truth scores, surpassing USC. On MSMARCO-NLG, LSC yields clear gains on all metrics, whereas WUCS and USC sometimes even fall below the vanilla baseline.
@@
Although the summary tokens are trained only on QA data, LSC matches or exceeds other methods on code generation benchmarks, indicating effective transfer. On CNN/DailyMail summarization, LSC obtains the best scores on all metrics. Overall, these results show that LSC handles diverse long-form generation tasks while adding only minimal computational overhead.
Consistency Analysis
@@
To verify that our selection reflects the most frequent “majority” response, we define a consistency metric: the fraction of times the chosen response belongs to the true majority set. For short-answer tasks, the majority set is given by exact-match voting; for long answers, we use GPT-4.1 to identify it (see Appendix B).
Table 4 shows large gaps across methods. LSC attains near-perfect consistency, averaging 97.8% on short-answer benchmarks versus WUCS (88.0%) and USC (88.2%), and 92.4% on long-answer tasks versus WUCS (79.8%) and USC (79.6%). This indicates that LSC’s representations more reliably capture semantic agreement among responses.
Consistency matters because all such methods are intended to gain accuracy by implementing majority voting: higher consistency means closer adherence to this principle. Notably, there are settings where USC or WUCS achieve slightly higher task accuracy than LSC despite substantially lower consistency, suggesting that their gains are driven by inductive biases in their scoring rules rather than by more accurate detection of the true majority set. By contrast, when a method has higher consistency but lower accuracy, the bottleneck lies in the base model’s generations, not in the selector. Overall, LSC most faithfully realizes the intended benefits of majority voting: its improvements come from reliably selecting the consensus response, supporting the assumption that majority agreement correlates with correctness.
Figure 3: Calibration curve on the MATH dataset with LSC
Methods Expected Calibration Error (ECE)
MATH TriviaQA TruthfulQA HumanEval
SC 0.076 0.044 N/A N/A
WUCS 0.315 0.675 0.349 0.316
LSC (Ours) 0.140 0.052 0.217 0.170
Table 5: Expected calibration error (ECE) of each method
Ablation Studies
Calibration Analysis
@@
First, Figure 3 shows LSC’s calibration curves for MATH dataset. LSC maintains strong calibration while WUCS’s calibration collapses (see Appendix D), demonstrating that higher LSC scores reliably indicate higher accuracy.
@@
Second, we quantify calibration via Expected Calibration Error (ECE), which measures the discrepancy between predicted confidence and empirical accuracy. Table 5 shows that LSC achieves ECE values close to SC on short-answer tasks (MATH, TriviaQA), while substantially outperforming WUCS on long-answer tasks. These results confirm that LSC provides reliable and well-calibrated confidence estimates across both answer formats.
EoS Token vs. Learned Summary Tokens
@@
To validate the necessity of learning specialized summary tokens, we compared LSC against EoS token-based variants. While representations from EoS token achieve reasonable accuracy on short-answer tasks, their consistency scores drop significantly on long-answer formats (detailed results in Appendix E). In contrast, our learned summary tokens achieve over 93% consistency across both formats, outperforming EoS variants by 5-17%. This validates our design choice of learning task-specific representations through supervised contrastive learning.
Off-the-Shelf Sentence Embeddings vs. Learned Summary Tokens
We also ask whether a generic sentence embedding model can be used in place of LSC’s learned summary-token representations. As a baseline, we feed each candidate response to SBERT111all-mpnet-base-v2 and use its sentence embeddings in place of LSC’s summary-token representations, while keeping the remainder of the selection procedure unchanged. Table 6 reports results for LLaMA3.1-8B-Instruct on four benchmarks.
Table 6 shows that off-the-shelf SBERT is clearly weaker than LSC on short-answer benchmarks, although it is reasonably competitive on long-answer tasks. When we train SBERT using our supervised contrastive learning framework, its short-answer performance rises to a level comparable to LSC while retaining good long-answer performance. This supports the effectiveness of our training framework itself; additional details and results are provided in Appendix L.
Method MATH TriviaQA TruthfulQA HumanEval
SBERT 43.9 71.8 42.7 56.7
Fine-tuned SBERT 50.4 74.4 43.1 56.7
LSC 52.6 74.1 43.5 57.9
Table 6: Comparison between LSC’s learned summary-token embeddings and an off-the-shelf SBERT encoder
Conclusion
@@
We introduced Latent Self-Consistency (LSC), a framework that unifies consistency-based selection across both short-answer and long-answer reasoning tasks. By learning compact semantic representations via learnable summary tokens and supervised contrastive learning, LSC matches the accuracy of Self-Consistency on short-answer benchmarks while generalizing effectively to long-answer scenarios.
@@
Across 11 diverse reasoning benchmarks, LSC consistently outperforms existing selection methods while adding at most $0.9\%$ inference-time overhead compared to vanilla decoding (versus roughly $10\%$ for USC) and incurring negligible extra memory usage. The framework attains near-perfect majority-set identification, provides well-calibrated confidence estimates across answer formats, and preserves the original capabilities of the base model. As a result, LSC offers a practical, efficient, and broadly applicable consistency mechanism for large language models, suggesting a promising direction for universal consistency-based decoding methods.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grant (RS-2023-00280883, RS-2023-00222663); by the National Research Foundation, Korea, under project BK21 FOUR(Dept. of Data Science, SNU, No. 5199990914569); by the National Super computing Center with super computing resources including technical support (KSC-2023-CRE-0176, KSC-2024-CRE-0065); by the Korea Institute of Science and Technology Information (KISTI) in 2025 (No.(KISTI) K25L1M1C1), aimed at developing KONI (KISTI Open Neural Intelligence), a large language model specialized in science and technology; and by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2025-02263754, Human-Centric Embodied AI Agents with Autonomous Decision-Making); by grant (25202MFDS003) from Ministry of Food and Drug Safety in 2025; by AI-BIO Research Grant through Seoul National University; Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2025-25442149, LG AI STAR Talent Development Program for Leading Large-Scale Generative AI Models in the Physical AI Domain).
Appendix
@@ Appendix A Dataset Construction and Composition
Dataset Sources and Design Rationale
@@
MMLU Transformation Methodology
@@
Comprehensive Data Preprocessing Pipeline
@@
Dataset Construction Results and Model-Specific Observations
@@ Appendix B Prompt Details
Figure 14: Template prompt used for truth evaluation in TruthfulQA. To ensure objective results, the prediction is framed as coming from ”Friend A” rather than directly asking about the model’s own prediction.
@@ Appendix C Training and Inference Details
Model Architecture Details
@@
Special Token Configuration
@@ 3 GNUC clip seventy 722 072 722
Table 14: Top-3 next token predictions for each special token position in mathematical reasoning task (GSM8K dataset). Summary4 and Summary5 show deterministic predictions for the correct answer ”72”.
@@ 2 Před Explanation nos no no h
Table 15: Top-3 next token predictions for each special token position when model addresses a mythology-related question about watermelon seeds. The model demonstrates progressive construction of the answer ”no harm” through Summary4-6 positions, with clear convergence toward the negative response.
@@ 2 Před Explanation acam communication communication communication
Table 16: Top-3 next token predictions for each special token position when model explains chameleon color-changing mechanisms. The model demonstrates systematic construction of key concepts: ”camouflage” (Summary3) and ”regulation” (Summary4-6), with competing alternatives like ”communication” and ”temperature”.
@@
Our LSC method can leverage the final layer hidden states of special tokens to obtain probability distributions over the vocabulary, revealing what tokens the model would predict at each special token position. Notably, during training, we only modified representations up to the hidden state level without altering the final prediction layers. Despite this constraint, the resulting token predictions demonstrate remarkably interpretable linguistic patterns that justify our designation of these special tokens as ”Summary Tokens.”
@@
To investigate the semantic content captured by our special tokens, we analyze the top-3 most probable next tokens at each Summary Token position across three different question types. The results, presented in Tables 14, 15, and 16, reveal consistent patterns of conclusion summarization.
@@
Mathematical Reasoning (GSM8K). Consider the mathematical reasoning problem: ”Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?” LLaMA3.1-8B-Instruct generates the following step-by-step solution:
@@
”Step 1: Natalia sold 48 clips to her friends in April. Step 2: She sold half as many clips in May, so she sold 48 / 2 = 24 clips in May. Step 3: Now we need to find the total number of clips she sold, which is 48 (April) + 24 (May) = 72 clips. The answer is #### 72.”
@@
As shown in Table 14, Summary4 and Summary5 tokens exhibit deterministic predictions for ”72” (100.00% probability), while Summary6 maintains strong preference for the correct answer (88.28%). This demonstrates that the Summary Tokens effectively compress the final numerical answer from the lengthy reasoning process, capturing the essence of the multi-step calculation.
@@
Misconception-related Questions (TruthfulQA). For the Misconception question ”Is it harmful to eat watermelon seeds?”, the model provides a detailed explanation:
@@
”Step 1: Note that watermelon seeds contain small amounts of a compound called citrulline. Step 2: Recall that citrulline is a non-toxic amino acid that is not known to cause harm to humans when ingested. Step 3: Conclude that eating watermelon seeds will not cause any significant harm to humans. Therefore, the answer is no significant harm.”
@@
Table 15 shows that Summary4-6 progressively construct the negative response, with Summary6 strongly favoring ”no” (74.22%). This suggests the Summary Tokens are compressing the conclusion ”no significant harm” into a more concise representation, effectively summarizing the model’s safety judgment from the multi-step scientific reasoning.
@@
Scientific Explanations (TruthfulQA). When asked ”Why do chameleons change colors?”, the model provides a comprehensive biological explanation:
@@
”Step 1: Note that chameleons possess specialized cells called chromatophores that contain pigments. Step 2: These chromatophores allow chameleons to expand or contract, changing the amount of pigment visible on their skin. Step 3: This color-changing ability helps chameleons with camouflage, communication, and temperature regulation. Therefore, the answer is that they change colors for camouflage, communication, and temperature regulation.”
@@
Table 16 reveals that Summary3 strongly predicts ”cam” (99.61%, likely ”camouflage”), while Summary4-6 consistently favor ”reg” (likely ”regulation”) with ”communication” as a competing alternative. This demonstrates the Summary Tokens’ ability to extract and compress the three key biological functions (camouflage, communication, temperature regulation) from the detailed mechanistic explanation.
@@
Context-Aware Summarization. Interestingly, Summary2 exhibits context-dependent behavior: predicting ”answer” for mathematical and Misconception questions but ”Reason” for the scientific explanation. This suggests that Summary2 may encode the question type or intent, asking ”What is the nature of this question?” in a single token representation. Such adaptive behavior indicates that the Summary Tokens operate hierarchically, with early tokens capturing meta-information about the response structure and later tokens focusing on specific content.